The rate of ice formation is (2- 0.3t) grams/minutes. If there was initial 10grams of ice the ice maker, find the mass of the ice after 10minutes
1
Expert's answer
2019-07-19T13:55:29-0400

Since rate of ice formation depends on time, the simple equation


"m={{m}_{in}}+v\cdot t"


where min is the initial ice mass, v is the rate of ice formation, is wrong.

The mass of the ice after t minutes can be calculated by the equation


"m={{m}_{in}}+\int\limits_{0}^{t}{v\left( \tau \right)d\tau }"


We are given "v(t)=(2-0.3t)"grams/minutes. Thus


"m={{m}_{in}}+\int\limits_{0}^{t}{\left( 2-0.3\tau \right)d\tau }"


Integrate


"m={{m}_{in}}+\left. \left( 2\tau -0.3\frac{{{\tau }^{2}}}{2} \right) \right|_{0}^{t} 

m={{m}_{in}}+2t-0.3\frac{{{t}^{2}}}{2}"

Substitute min =10grams and t=10minutes


"m=10+2\cdot 10-0.3\frac{{{10}^{2}}}{2}=10+20-15=15\,grams"

Thus, the mass of ice after 10 minutes is equal to 15 grams